# Toward regular black holes in sixth-derivative gravity

### v1

@@ Abstract
We study spherically symmetric static solutions of the most general sixth-derivative gravity using series expansions. Specifically, we prove that the only solutions of the complete theory (i.e., with generic coupling constants) that possess a Frobenius expansion around the origin, $r=0$, are necessarily regular. When restricted to specific branches of theories (i.e., imposing particular constraints on the coupling constants), families of potentially singular solutions emerge. By expanding around $r=r_{0}\neq 0$, we identify solutions with black-hole horizons. Finally, we argue that, unlike in fourth-derivative gravity, the conditions $R=0$ and $g_{tt}g_{rr}=-1$ are too restrictive for sixth-derivative gravity solutions.
@@ I Introduction
In the last decade there has been a renewed interest in the important yet ambiguous role of higher derivatives in gravity. The perturbative quantization of Einstein gravity and even the quantization of matter fields on a curved background require the introduction of higher-derivative terms in the gravitational sector to renormalize loop divergences [1]. Such terms also appear in the low-energy regime of string theory [2, 3]. Alternatively, by considering higher-derivative terms at fundamental level one can formulate renormalizable [4] and super-renormalizable [5] models of quantum gravity. This comes with the drawback of having ghost-like particles in the spectrum, traditionally associated with violation of unitarity and other instabilities. Recently, however, considerable effort has been made to analyze scenarios in which the effect of ghosts can be controlled (at classical and quantum levels) and unitarity can be recovered [6, 7, 8, 9, 10, 11]. Some of these constructions depend on more complicated higher-derivative actions, with metric derivatives higher than fourth [11] or even in the form of non-local operators [12, 13, 14, 15].
@@
Among the main issues that a quantum theory of gravity is expected to address it is the resolution of the spacetime singularities that occur in general relativity. In the absence of a satisfactory theory and an underlying regularization mechanism, phenomenological models of quantum-corrected regular black holes abound in the literature. The question of whether higher derivatives or non-locality could solve the problem already at classical level has been considered in different contexts [14, 15, 16, 17, 15, 18, 21, 22, 19, 20], but owing to the complicated equations of motion and the humungous amount of possible higher-derivative correction terms, such studies most often involve a linearization of field equations. In this simplified setup it was shown that there is a significant difference between renormalizable theories defined by actions with four and with more than four metric derivatives. For instance, the latter models have a regular non-relativistic limit when coupled to normal matter, while the former still displays curvature singularities [19, 20].
@@
Nevertheless, results concerning exact solutions of higher-derivative gravities are nearly absent in the literature. An important exception is the case of fourth-derivative gravity, for which several families of exact solutions are known (see, e.g.[23, 24, 25, 26, 27, 28, 29, 36, 30, 31, 34, 32, 35, 33, 37, 38] and references therein). It turns out that this model might not offer a resolution of the singularity problem at classical level. Indeed, it admits singular static spherically symmetric solutions — and the asymptotically flat solutions that couple to normal ghost-free matter appear to contain a naked singularity [26].
@@
A suggestion that the situation could be different for gravity models with more derivatives was provided in the work [24], but a systematic study of the models and their solutions remained open. Recently, some exact solutions for the Einstein gravity augmented by the six-derivative Goroff–Sagnotti term have been obtained [39]. While this term can be regarded as a two-loop correction based on the quantization of general relativity [40], it is not the only one at this perturbative order. Moreover, taken alone it does not shed light to the features of the solutions of super-renormalizable gravity models, nor does it offer an immediate insight on the singularity problem, as divergent solutions also seem to exist [41].
@@
Our goal in this short paper is to present some results regarding the static spherically symmetric solutions in gravity models defined by actions with up to six derivatives of the metric, especially concerning the occurrence of regular spacetime configurations. In the same spirit as [26], our basic assumption is that such higher-derivative action is the relevant one at some energy scale, regardless of whether it is a fundamental theory or an emergent one. We shall not attempt to discuss the complicated problem of ghosts and stability of the solutions, as this lies beyond the scope of this work.
II General sixth-derivative gravity
@@
$S$ $=$ $\int\mathrm{d}^{4}x\sqrt{-g}\Bigg{[}\alpha R+\beta_{1}R^{2}+\beta_{2}R_{\mu\nu}^{2}+\gamma_{1}R\Box R+\gamma_{2}R_{\mu\nu}\Box R^{\mu\nu}+\gamma_{3}R^{3}+\gamma_{4}RR_{\mu\nu}R^{\mu\nu}+\gamma_{5}R_{\mu\nu}R^{\mu}{}_{\rho}R^{\rho\nu}$ (1)
$+\,\gamma_{6}R_{\mu\nu}R_{\rho\sigma}R^{\mu\rho\nu\sigma}+\gamma_{7}RR_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}+\gamma_{8}R_{\mu\nu}R^{\mu}{}_{\rho\sigma\tau}R^{\nu\rho\sigma\tau}+\gamma_{9}R_{\mu\nu\rho\sigma}R^{\mu\nu\tau\upsilon}R^{\rho\sigma}{}_{\tau\upsilon}+\gamma_{10}R_{\mu\rho\nu\sigma}R^{\mu}{}_{\tau}{}^{\nu}{}_{\upsilon}R^{\rho\tau\sigma\upsilon}\Bigg{]},$
where the constants $\alpha$, $\beta_{1,2}$ and $\gamma_{1,\ldots,10}$ are, respectively, the coefficients of the terms with a total number of 2, 4 and 6 derivatives. Any other four- or six-derivative term can be cast as a combination of the terms in (1) and boundary or topological terms, that do not contribute to the equations of motion [42].
The variation of the above action with respect to the metric $g_{\mu\nu}$ yields the equations of motion $H_{\mu\nu}=0$, which we calculated using the package xAct [43, 44] for Mathematica [45]. For a generic metric in the standard spherically symmetric coordinates,
@@
the field equations assume a diagonal form. Moreover, the generalized Bianchi identity $\nabla^{\nu}H_{\mu\nu}=0$ acts as a constraint and we end up with only two independent equations, that can be taken to be $H_{tt}=0$ and $H_{rr}=0$. Together they form a system of coupled differential equations that is of sixth order for $B(r)$ and fifth order for $A(r)$. The terms of highest differential order are originated from the structures proportional to $\gamma_{1}$ and $\gamma_{2}$ in (1), as these are the ones that contain the largest number of derivatives acting on a single metric component.
III Solutions of the general sixth-derivative gravity
Solutions of the system around a certain point $r=r_{0}$ can be obtained by assuming that the functions $A(r)$ and $B(r)$ are represented by Frobenius series. For expansions around $r=0$ we use the ansatz
$A(r)=r^{s}\sum_{n=0}^{\infty}a_{n}r^{n},\qquad B(r)=b_{0}r^{t}\left(1+\sum_{n=1}^{\infty}b_{n}r^{n}\right),$ (3)
@@ III Solutions of the general sixth-derivative gravity
$F(r)=\Delta^{w}\sum_{n=0}^{\infty}f_{n}\Delta^{n},\quad B(r)=b_{0}\Delta^{t}\left(1+\sum_{n=1}^{\infty}b_{n}\Delta^{n}\right),\quad\Delta\equiv r-r_{0}$ (4)
@@
with $w,t\in\mathbb{R}$ to be determined. The corresponding expansion of the field equations at lowest order constitute the system of indicial equations, for they determine the admissible values of $s$ (or $w$) and $t$. After fixing this pair of parameters one can, in principle, solve the field equations order by order. We refer to the possible families of solutions by the pair of indexes $(s,t)_{0}$ and $(w,t)_{r_{0}}$, with the subscript label indicating whether this is a solution around $r=0$ or $r=r_{0}\neq 0$.
@@
From the differential order of the field equations, one might expect that solutions could have up to 11 free parameters [this counting includes the parameter $b_{0}$, which is not physical for it corresponds to the time re-scaling freedom of the metric (2)]. Although the field equations form a nonlinear system, the reasoning based on the differential order seems to work in quadratic gravity [26]. As we show below, all the solutions we found satisfy this upper bound.
III.1 Solutions around $r=0$
If all the coefficients in the action (1) are non-zero and completely arbitrary (i.e., no relation between them is assumed), the system of indicial equations around $r=0$ only admit the solution
$s=t=0.$ (5)
This result is obtained by calculating the field equations and the corresponding indicial equations for each of the terms in the action (1). The assumption that the coefficients in the action are completely arbitrary means that all those indicial equations must be satisfied at the same time and it turns out that the solution $s=t=0$ is the only one common to all of them. In other words, there is only one family of solutions of the type (3) around $r=0$, with indicial structure $(0,0)_{0}$.
@@ III.1 Solutions around $r=0$
After solving the equations of $H_{tt}=0$ and $H_{rr}=0$ up to nine orders in $r$, we are convinced that the solutions in this family are characterized by six parameters, which can be taken to be $a_{2}$, $a_{3}$, $a_{4}$, $b_{0}$, $b_{2}$ and $b_{4}$. Among these, only five parameters are physical, for $b_{0}$ corresponds to the time re-scaling freedom of the metric. The free parameters appeared within the first five orders of the expansion; beyond this, at each new order we found two equations for two new parameters. The general structure of the solution is
$A(r)$ $=$ $1+a_{2}r^{2}+a_{3}r^{3}+a_{4}r^{4}+\frac{b_{3}\bar{a}_{5}}{\gamma_{2}(3\gamma_{1}+\gamma_{2})}r^{5}+O(r^{6}),$
$\frac{B(r)}{b_{0}}$ $=$ $1+b_{2}r^{2}+b_{3}r^{3}+b_{4}r^{4}+\frac{b_{3}\bar{b}_{5}}{\gamma_{2}(3\gamma_{1}+\gamma_{2})}r^{5}+O(r^{6}),$
@@
$3(4\gamma_{1}+\gamma_{2})b_{3}=(8\gamma_{1}+3\gamma_{2})a_{3},$ (7)
@@
and the quantities $\bar{a}_{5}$ and $\bar{b}_{5}$ depend polynomially on the coefficients $a_{2}$ and $b_{2}$, and on the parameters of the model. The coefficients of the terms $O(r^{6})$ are determined from the lower order ones.
@@
A remarkable feature of these solutions is that the geometry is regular at $r=0$ in the sense that all the curvature invariants constructed by contracting an arbitrary number of Riemann and metric tensors are bounded. This happens because the solution in Eq. (LABEL:Sol00Geral) satisfies $a_{0}=1$ and $a_{1}=b_{1}=0$. In particular, for the Kretschmann scalar we have
$R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}\underset{r\to 0}{=}12\left(a_{2}^{2}+b_{2}^{2}\right)+O(r).$ (8)
@@
It is also interesting to notice that the solution in Eq. (LABEL:Sol00Geral) explicitly requires $\gamma_{2}(3\gamma_{1}+\gamma_{2})\neq 0$. This is precisely the condition for the model (1) to have sixth-order derivatives in both the spin-2 and spin-0 sectors. If this condition is not satisfied, there might still be non-trivial solutions of the type $(0,0)_{0}$, but the structure and characteristic number of parameters of the solutions can be different from the ones of the most general model. For instance, if $\gamma_{1}=\gamma_{2}=0$ the field equations contain at most four derivatives acting on a single metric function, instead of six, and one might expect a reduction of the number of free parameters in a solution. In Sec. IV we shall discuss some important cases of incomplete sixth-derivative models, i.e., in which some coefficients $\gamma_{i}$ are null or assume other particular values.
III.2 Solutions around $r=r_{0}\neq 0$
@@
Differently from the solutions around $r=0$, we have identified several families of solutions around a finite point $r=r_{0}\neq 0$. This diversity is indicative of the various possibilities for the point $r_{0}$, which can be, e.g., a horizon or a generic point, depending on the indicial structure of the solution (4).
@@
The solutions around a generic point are in the class $(0,0)_{r_{0}}$. After solving the field equations up to three orders in $r-r_{0}$, we are convinced that such solutions are characterized by 11 free parameters ($f_{0},\ldots,f_{4}$, $b_{0},\ldots,b_{4}$ and $r_{0}$), which appear already at first order. This is precisely the maximal number of free parameters (including the non-physical parameter $b_{0}$) that we expected for a solution of the type (4), taking into account the differential order of the field equations. The situation here is analogous to the fourth-derivative gravity, where the family of solutions $(0,0)_{r_{0}}$ around a generic point also has the maximal number of free parameters [26]. This counting of free parameters is confirmed by the analysis using the metric written in conformal-to-Kundt form, whose details will be reported elsewhere [41].
@@
Expansions around a horizon $r=r_{0}$ correspond to the family $(1,1)_{r_{0}}$. Similar analysis suggests that these solutions are characterized by 6 physical parameters and their existence allows us to conclude that there are black holes in sixth-derivative gravity. Since asymptotically flat regular black holes must have an even number of horizons, one expects to also encounter families of solutions expanded around a double (extreme) horizon. The search for such solutions is more complicated precisely because of their extremal nature, which makes them to be completely determined by the parameters of the model (with the only free parameters being the gauge one). Again, more details about these and other solutions will be provided in a separate work [41].
III.3 Solutions with $R=0$ and with $g_{tt}g_{rr}=-1$

### v2

@@ Abstract
We study spherically symmetric static solutions of the most general sixth-derivative gravity using series expansions. Specifically, we prove that the only solutions of the complete theory (i.e., with generic coupling constants) that possess a Frobenius expansion around the origin, $r=0$, are necessarily regular. When restricted to specific branches of theories (i.e., imposing particular constraints on the coupling constants), families of potentially singular solutions emerge. By expanding around $r=r_{0}\neq 0$, we identify solutions with black hole horizons. Finally, we argue that, unlike in fourth-derivative gravity, the conditions $R=0$ and $g_{tt}g_{rr}=-1$ are too restrictive for sixth-derivative gravity solutions.
@@ I Introduction
In the last decade there has been a renewed interest in the important yet ambiguous role of higher derivatives in gravity. The perturbative quantization of Einstein gravity and even the quantization of matter fields on a curved background require the introduction of higher-derivative terms in the gravitational sector to renormalize loop divergences [1, 2, 3]. Such terms also appear in the low-energy regime of string theory [4, 5]. Alternatively, by considering higher-derivative terms at the fundamental level, one can formulate renormalizable and superrenormalizable models of quantum gravity [6, 7]. This comes with the drawback of having ghostlike particles in the spectrum, traditionally associated with violation of unitarity and other instabilities. Recently, however, considerable effort has been made to analyze scenarios in which the effect of ghosts can be controlled (at classical and quantum levels) and unitarity can be recovered [8, 9, 10, 11, 12, 13]. Some of these constructions depend on more complicated actions, with metric derivatives higher than fourth [7, 13] or in the form of nonlocal operators [14, 15, 16, 17].
@@
Among the main issues that a quantum theory of gravity is expected to address is the resolution of the spacetime singularities that occur in general relativity. In the absence of a satisfactory theory and an underlying regularization mechanism, phenomenological models of quantum-corrected regular black holes abound in the literature. The question of whether higher derivatives could solve the problem has been considered in different contexts [16, 17, 18, 19, 17, 20, 25, 24, 23, 26, 21, 22], but owing to the complicated equations of motion and the humongous amount of possible higher-derivative correction terms, such studies most often involve a linearization of field equations. In this simplified setup it was shown that there is a significant difference between theories defined by actions with four and with more than four metric derivatives. For instance, the latter models have a regular nonrelativistic limit when coupled to normal matter, while the former still displays curvature singularities [21, 22].
@@
Nevertheless, results concerning exact solutions of higher-derivative gravities are nearly absent in the literature. An important exception is the case of fourth-derivative gravity, for which several families of exact solutions are known (see, e.g., [27, 28, 29, 30, 31, 32, 33, 34, 41, 35, 36, 39, 37, 40, 38, 42, 43] and references therein). It turns out that this model might not offer a resolution of the singularity problem at the classical level. Indeed, it admits singular static spherically symmetric solutions — and the asymptotically flat solutions that couple to normal ghost-free matter appear to contain a naked singularity [31].
@@
A suggestion that the situation could be different for gravity models with more derivatives was provided in the work [28], which reported finding regular solutions, but a systematic study of the models and their solutions remained open. Recently, exact solutions for the Einstein gravity augmented by the six-derivative term ${C^{\mu\nu}}_{\alpha\beta}{C^{\alpha\beta}}_{\rho\sigma}{C^{\rho\sigma}}_{\mu\nu}$ have been obtained [44]. While this term can be regarded as a two-loop correction based on the quantization of general relativity [3], it is not the only one at this perturbative order. Moreover, taken alone it does not shed light on the features of the solutions of superrenormalizable gravity models, nor does it offer an immediate insight into the singularity problem, as divergent solutions also seem to exist [45].
Exact solutions are also known for the Einsteinian cubic gravity and higher-dimensional quasitopological theories of gravity [47, 48, 49, 49, 46]. Both classes of models are based on actions with very specific combinations of higher-derivative terms in order to yield second-order field equations for static spherically symmetric metrics. However, owing to the uniqueness of such constructions, their solutions ought to be the exception rather than the rule for higher-derivative gravities. For instance, regular black holes have been recently obtained in this framework but only in higher-dimensional theories with an infinite tower of higher-derivative terms [50, 51].
@@
Our goal in this short paper is to present results regarding the static spherically symmetric vacuum solutions in gravity models defined by actions with up to six derivatives of the metric, especially concerning the occurrence of regular spacetime configurations. The basic assumption is that such higher-derivative action is the relevant one at some energy scale, regardless of whether it is a fundamental theory or an emergent one. We shall not attempt to discuss the complicated problem of ghosts and stability of the solutions, as this lies beyond the scope of this work.
@@
$S$ $=$ $\int\mathrm{d}^{4}x\sqrt{-g}\Big{[}\alpha R+\beta_{1}R^{2}+\beta_{2}R_{\mu\nu}^{2}+\gamma_{1}R\Box R+\gamma_{2}R_{\mu\nu}\Box R^{\mu\nu}$ (1)
$+\,\gamma_{3}R^{3}+\gamma_{4}RR_{\mu\nu}R^{\mu\nu}+\gamma_{5}R_{\mu\nu}R^{\mu}{}_{\rho}R^{\rho\nu}+\gamma_{6}R_{\mu\nu}R_{\rho\sigma}R^{\mu\rho\nu\sigma}$
$+\,\gamma_{7}RR_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}+\,\gamma_{8}R_{\mu\nu\rho\sigma}R^{\mu\nu\tau\upsilon}R^{\rho\sigma}{}_{\tau\upsilon}\Big{]},$
where the constants $\alpha$, $\beta_{1,2}$, and $\gamma_{1,\ldots,8}$ are, respectively, the coefficients of the terms with a total number of 2, 4, and 6 derivatives. Any other four- or six-derivative term can be cast as a combination of the terms in (1) and boundary or topological terms (that do not contribute to the equations of motion) [53, 52].
The variation of the above action with respect to the metric yields the field equations
$H_{\mu\nu}\equiv\frac{1}{\sqrt{-g}}\frac{\delta S}{\delta g^{\mu\nu}}=0,$ (2)
which we calculated using the package xAct [54, 55, 56] for Mathematica [57]. For a generic metric in the standard spherically symmetric coordinates,
@@
the field equations assume a diagonal form. Moreover, the generalized Bianchi identity $\nabla^{\nu}H_{\mu\nu}=0$ acts as a constraint, and we end up with only two independent equations that can be taken to be $H_{tt}=0$ and $H_{rr}=0$. Together, they form a system of coupled differential equations that is of sixth order for $B(r)$ and fifth order for $A(r)$. The terms of highest differential order are originated from the structures proportional to $\gamma_{1}$ and $\gamma_{2}$ in (1), as these are the ones that contain the largest number of derivatives acting on a single metric component.
Solutions of the field equations around a certain point $r=r_{0}$ can be obtained by assuming that the functions $A(r)$ and $B(r)$ are represented by Frobenius series. For expansions around $r=0$, we use the ansatz
$A(r)=r^{s}\sum_{n=0}^{\infty}a_{n}r^{n},\qquad\frac{B(r)}{b_{0}}=r^{t}\left(1+\sum_{n=1}^{\infty}b_{n}r^{n}\right),$ (4)
@@
$F(r)=\Delta^{w}\sum_{n=0}^{\infty}f_{n}\Delta^{n},\quad\frac{B(r)}{b_{0}}=\Delta^{t}\left(1+\sum_{n=1}^{\infty}b_{n}\Delta^{n}\right),\quad\Delta\equiv r-r_{0}$ (5)
@@
with $w,t\in\mathbb{R}$ to be determined. We assume that $a_{0},b_{0},f_{0}\neq 0$, so that the powers $s$, $t$, and $w$ define the leading terms of the series. The corresponding expansion of the field equations at the lowest order comprises the system of indicial equations, for they determine the admissible values of $s$ (or $w$) and $t$. This is a first constraint on the space of possible solutions. After fixing this pair of parameters, a solution might be obtained by solving the field equations order by order. We refer to the families of solutions by the pair of indexes $(s,t)_{0}$ and $(w,t)_{r_{0}}$, with the subscript label indicating whether this is a solution around $r=0$ or $r=r_{0}\neq 0$.
@@
From the differential order of the field equations, one might expect that solutions could have up to 11 free parameters [this counting includes the parameter $b_{0}$, which is not physical for it corresponds to the time rescaling freedom of the metric (3)]. Although the field equations form a nonlinear system, the reasoning based on the differential order seems to work in quadratic gravity [31]. As we show below, all the solutions we found satisfy this upper bound.
II Solutions around $r=0$
The expanded field equations at the lowest order using the ansatz (4) have the general structure
$\begin{split}\frac{H_{tt}(r)}{B(r)}&=r^{p(s)}\sum_{i=1}^{8}\gamma_{i}\,g_{i}(s,t,a_{0})+\ldots=0,\\ H_{rr}(r)&=r^{q(s)}\sum_{i=1}^{8}\gamma_{i}\,h_{i}(s,t,a_{0})+\ldots=0,\end{split}$ (6)
where the powers $p$ and $q$ depend on $s$, the coefficients $g_{i}$ and $h_{i}$ only depend on $s$, $t$, and $a_{0}$, and the ellipsis denote terms of higher order in $r$. The explicit expressions for all these quantities are presented in the Appendix A. Notice that, to the lowest order, (6) only receives contributions from the six-derivative terms in the action (1).
The requirement that the field equations are solved for arbitrary values of the parameters $\gamma_{1,\ldots,8}$, results in the system of indicial equations
$g_{i}(s,t,a_{0})=h_{i}(s,t,a_{0})=0,\quad\forall\,i=1,\ldots,8.$ (7)
A detailed consideration of this system is carried out in the Appendix A, where we also prove that its only solution is
$s=t=0,\qquad a_{0}=1.$ (8)
In other words, there is only one family of solutions of the type (4) around $r=0$, with indicial structure $(0,0)_{0}$.
@@ II Solutions around $r=0$
After solving the equations $H_{tt}=0$ and $H_{rr}=0$ up to nine orders in $r$, we are convinced that the solutions in this family are characterized by six parameters, which can be taken to be $a_{2}$, $a_{3}$ (or $b_{3}$), $a_{4}$, $b_{0}$, $b_{2}$, and $b_{4}$. Among these, only five parameters are physical, for $b_{0}$ corresponds to the time rescaling freedom of the metric. The free parameters appeared within the first five orders of the expansion; beyond this, at each new order we found two equations for two new parameters. The general structure of the solution is
$A(r)$ $=$ $1+a_{2}r^{2}+a_{3}r^{3}+a_{4}r^{4}+\frac{(a_{3}+b_{3})\bar{a}_{5}}{\gamma_{2}(3\gamma_{1}+\gamma_{2})}r^{5}+O(r^{6}),$
$\frac{B(r)}{b_{0}}$ $=$ $1+b_{2}r^{2}+b_{3}r^{3}+b_{4}r^{4}+\frac{(a_{3}+b_{3})\bar{b}_{5}}{\gamma_{2}(3\gamma_{1}+\gamma_{2})}r^{5}+O(r^{6}),$
@@
$(8\gamma_{1}+3\gamma_{2})a_{3}=3(4\gamma_{1}+\gamma_{2})b_{3},$ (10)
@@
and the quantities $\bar{a}_{5}$ and $\bar{b}_{5}$ depend polynomially on $a_{2}$ and $b_{2}$ and on the parameters of the model. The coefficients of the terms $O(r^{6})$ are determined from the lower-order ones.
@@
A remarkable feature of these solutions is that the geometry is regular at $r=0$ in the sense that all the curvature invariants constructed by contracting an arbitrary number of Riemann and metric tensors are bounded. This happens because the solution in Eq. (LABEL:Sol00Geral) satisfies $a_{0}=1$ and $a_{1}=b_{1}=0$. In particular, for the Kretschmann and Ricci scalars we have
$R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}$ =$\underset{r\to 0}{=}12\left(a_{2}^{2}+b_{2}^{2}\right)+O(r),$ (11)
$R$ =$\underset{r\to 0}{=}6\left(a_{2}-b_{2}\right)+O(r).$ (12)
@@
It is also interesting to notice that the solution (LABEL:Sol00Geral) explicitly requires $\gamma_{2}(3\gamma_{1}+\gamma_{2})\neq 0$. This is precisely the condition for the model (1) to have sixth-order derivatives in both the spin-2 and spin-0 sectors. If this condition is not satisfied, the structure of the solution family $(0,0)_{0}$ can be different from the one of the most general model. For instance, if $\gamma_{1}=\gamma_{2}=0$ the field equations contain at most four derivatives acting on a single metric function, and one might expect a reduction of the number of free parameters in a solution. In Sec. V we shall discuss some important cases of incomplete sixth-derivative models, i.e., in which some coefficients $\gamma_{i}$ are null or assume other particular values.
III Solutions around $r=r_{0}\neq 0$
@@
In contrast to the solutions around $r=0$, we have identified several families of solutions around a finite point $r=r_{0}\neq 0$. This diversity is indicative of the various possibilities for the point $r_{0}$, which can be, e.g., a horizon or a generic point, depending on the indicial structure of the solution (5).
@@
The solutions around a generic point are in the class $(0,0)_{r_{0}}$. After solving the field equations up to three orders in $r-r_{0}$, we are convinced that such solutions are characterized by 11 free parameters ($f_{0},\ldots,f_{4}$, $b_{0},\ldots,b_{4}$, and $r_{0}$), which appear already at first order. This is precisely the maximal number of free parameters (including the nonphysical parameter $b_{0}$) that we expected for a solution of the type (5), taking into account the differential order of the field equations. The situation here is analogous to the fourth-derivative gravity, where the family of solutions $(0,0)_{r_{0}}$ around a generic point also has the maximal number of free parameters [31]. This counting of free parameters is confirmed by the analysis using the metric written in conformal-to-Kundt form [45].
Expansions around a horizon $r=r_{0}$ correspond to the family $(1,1)_{r_{0}}$. Similar analysis reveals that these solutions are characterized by 5 physical parameters (which can be taken to be $f_{0}$, $f_{1}$, $b_{1}$, $b_{2}$, and $r_{0}$) and their existence allows us to conclude that there are black holes in sixth-derivative gravity. The general structure of the solution is
$F(r)$ $=$ $f_{0}\Delta+f_{1}\Delta^{2}+f_{2}\Delta^{3}+f_{3}\Delta^{4}+O(\Delta^{5}),$
$B(r)$ $=$ $b_{0}\left(\Delta+b_{1}\Delta^{2}+b_{2}\Delta^{3}+b_{3}\Delta^{4}\right)+O(\Delta^{5}),$ (13)
where four parameters among $f_{0}$, $f_{1}$, $f_{2}$, $b_{1}$, $b_{2}$ are independent, $b_{0}$ represents the residual gauge freedom, and the subsequent parameters $f_{n}$ and $b_{n}$ for $n\in\{3,4,\ldots\}$ can be expressed as functions of those and of $r_{0}$. Different from the solution $(0,0)_{0}$, in this case we refrain from displaying the structure of the first terms of the solution because the expressions are considerably longer. Nevertheless, Eq. (13) is enough to verify that the solutions of type $(1,1)_{r_{0}}$ have regular curvature scalars at the horizon $r=r_{0}$. In fact, the Kretschmann and Ricci scalars are bounded,
$R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}$ =$\underset{r\to r_{0}}{=}\frac{16+16f_{0}^{2}r_{0}^{2}+(3b_{1}f_{0}+f_{1})^{2}r_{0}^{4}}{4r_{0}^{4}}+O(\Delta),$ (14)
$R$ =$\underset{r\to r_{0}}{=}\frac{4-[f_{0}(8+3b_{1}r_{0})+f_{1}r_{0}]r_{0}}{2r_{0}^{2}}+O(\Delta).$ (15)
@@
Since asymptotically flat regular black holes must have an even number of horizons, one expects also to encounter families of solutions expanded around a double (extreme) horizon. The search for such solutions is more complicated precisely because of their extremal nature, which makes them completely determined by the parameters of the model (with the only free parameters being the gauge one). More details about these and other solutions around $r=r_{0}\neq 0$ will be provided in a separate work [45].
IV Solutions with $R=0$ and with $g_{tt}g_{rr}=-1$